# Alpha Berkeley: A Scalable Framework for the Orchestration of Agentic Systems

### v1

@@ Abstract
Coordinating workflows across heterogeneous control systems remains a central challenge in safety-critical environments such as scientific facilities, industrial plants, and energy infrastructures. Language-model–driven agents offer a natural interface for these tasks, but existing approaches often lack scalability, reliability, and human oversight. We introduce the Alpha Berkeley Framework, a production-ready architecture for scalable agentic systems that integrate conversational context with robust tool orchestration. The framework features dynamic capability classification to select only relevant tools per task, a plan-first orchestration model that generates execution plans with explicit dependencies and optional human approval, context-aware task extraction that combines dialogue history with external memory and domain resources, and production-ready execution environments with checkpointing, artifact management, and modular deployment. We demonstrate its versatility through two case studies: a tutorial-style wind farm monitoring example and a deployment at the Advanced Light Source particle accelerator. These results establish Alpha Berkeley as a reliable and transparent framework for agentic systems in high-stakes domains.
Figure 1: Overview of the Alpha Berkeley Framework workflow. Multi-turn conversational input and external data sources are first processed into a structured task. Relevant capabilities are dynamically classified on a turn-by-turn basis, and the selected tools are passed to the execution planner. The planner generates a complete, inspectable execution plan with explicit dependencies, which is then carried out by the agent with context tracking and artifact management.
@@ I Introduction
Coordinating large-scale, safety-critical systems, such as smart electrical grids [1], industrial chemical processes [2], and scientific facilities such as particle accelerators [3], remains a central challenge in modern engineering. These environments integrate legacy and modern control systems that generate vast amounts of diverse data. Operators must act on this information under tight time constraints (for example, during accelerator fault recovery or power-grid disturbances) [4] yet retrieving, interpreting, and coordinating across systems remains cumbersome. Critical expertise is often preserved only as tacit human experience, scattered across institutional memory rather than codified in software. This tension between the need for flexibility and the rigidity of existing software architectures limits the ability to respond effectively to evolving operational demands and unexpected events.
Agentic AI [5], Language Model (LM) powered agents that orchestrate multi-step workflows through tool use, memory, and planning, offers a compelling opportunity to address these challenges. By providing a natural interface to data sources and control systems, LMs allow operators to express intent in high-level terms while delegating the orchestration of complex workflows to agents. In familiar domains such as web search [6] or travel planning [7], LMs already demonstrate this capability, as their training reflects widespread usage patterns and benchmark tasks. By contrast, in highly specialized domains such as particle accelerators, materials science beamlines, or other expert-driven facilities, much of the relevant expertise is absent from training corpora and cannot be inferred from general usage data. Instead, the knowledge is fragmented, highly context-dependent, and often confined to experts. Leveraging agentic AI in these environments therefore requires not only robust tool integration but also architectures capable of surfacing, formalizing, and operationalizing tacit human knowledge, while maintaining human-in-the-loop mechanisms to ensure safety and trust in high-stakes contexts.
In this paper, we introduce the Alpha Berkeley Framework [8], a production-ready architecture for building scalable agentic systems in complex scientific and industrial environments. Built on LangGraph [9] for its underlying graph-based orchestration substrate, our framework is distinguished by four characteristics:
@@
  • Production-ready adaptability: Designed for real deployments, the framework provides modular execution environments, artifact management, checkpointing, and diverse interfaces, including CLI, Open WebUI, and containerized services.
Together, these contributions move beyond proof-of-concept demonstrations toward a framework that is both flexible enough to adapt to new domains and robust enough for use in safety-critical operations. We illustrate the framework with two examples: a tutorial-style mock-up in a wind farm scenario, and a real-world deployment in the Advanced Light Source (ALS) control room. These case studies show both how the framework operates in practice and how it can be applied in complex, high-stakes environments.
@@ II.1 Agentic Systems and Tool Use
The foundations of agentic language models and multi-agent systems build on the classic definition of agents as entities that perceive their environment and act upon it, with rational agents aiming to maximize performance measures [10]. Extending this principle, multi-agent systems introduce shared environments in which agents may cooperate or compete. Early frameworks such as ReAct unified reasoning and acting in iterative loops for tool-based tasks [11], while AutoGen extended this paradigm toward multi-agent conversation orchestration, enabling specialized agents to collaborate on complex objectives [12].
Scaling tool use in LMs has since become a central concern. Toolformer demonstrated that LMs can be trained to make self-supervised API calls [13], while more recent work such as RAG-MCP addresses prompt bloat in tool selection by leveraging retrieval-augmented generation to select only relevant capabilities at inference time [14]. Beyond single-turn execution, memory and retrieval play an equally important role. Retrieval-Augmented Generation (RAG) integrates external knowledge into LM reasoning [15], but remains subject to context window limitations, such as the “Lost in the Middle” phenomenon [16]. To overcome this, memory-augmented agents like MemGPT [17] and subsequent surveys [18] have highlighted methods for extending long-context reasoning. Other work formalizes agent memory as a first-class capability, emphasizing the ability to retain past interactions to inform future steps [5]. Recent examples include SciBORG [19], a modular agentic framework for scientific tasks that employs FSA-based memory and context awareness to provide reliable execution and interpretable state transitions. Parallel efforts in task-oriented dialogue also offer complementary insights into grounded task extraction and intent modeling [20].
@@ II.2 Planning and Orchestration
Planning and orchestration in AI systems trace back to classical approaches such as STRIPS [21], PDDL [22], and hierarchical task network (HTN) planning [23]. Recent LM-based approaches build on these ideas by incorporating structured reasoning directly into language models, exemplified by Chain-of-Thought prompting [24] and Pre-Act, which combines multi-step planning with explicit reasoning steps [25]. DSPy further systematizes this direction, offering a modular framework for program synthesis and prompt optimization that reduces reliance on brittle, handcrafted prompt engineering [26, 27].
Complementing symbolic planning, graph-based architectures have emerged as powerful orchestration substrates for agents. LangGraph provides a task graph engine that enables explicit modeling of execution dependencies [9], while frameworks like PydanticAI contribute schema validation for structured outputs [28]. Related libraries, such as Instructor [29] and Outlines [30], offer constrained generation techniques that increase reliability in tool invocation.
Safety and human oversight remain critical in applying these systems to scientific and industrial domains. Reviews of self-driving laboratories emphasize that autonomy must integrate human intelligence for steering and safety [31], while approaches such as human-in-the-loop Bayesian experiment planning show how structured oversight can be embedded into experimental design [32]. More broadly, critical domains such as energy systems, particle accelerators, and laboratories require workflows with explicit approval points and safety boundaries.
@@ II.3 Domain-Specific Applications
Agentic AI has seen rapid adoption in scientific and industrial applications, where domain-specific requirements demand specialized integrations. In the natural sciences, autonomous laboratories have been identified as a key driver of accelerated discovery [33], supported by robotic labs and LM-driven pipelines for materials exploration [34]. Within imaging sciences, the PEAR framework has demonstrated multi-agent LM automation for ptychography workflows [35], while other efforts have highlighted emergent autonomous research capabilities of LMs across chemistry and materials science [36]. Several domain-focused systems illustrate this diversity: ChemCrow augments LMs with chemistry-specific tools [37], Co-scientist enables autonomous experimental planning in chemistry [38], and CRISPR-GPT applies agentic orchestration to gene-editing workflows [39].
Among large-scale scientific facilities, agentic approaches have also been explored for particle accelerators and synchrotron beamlines. GAIA demonstrated an early accelerator operations assistant combining retrieval, scripting, and control [40], while VISION developed a modular AI assistant for beamline experiments, focusing on natural human–instrument interaction [41]. GAIA and VISION illustrate how agentic approaches can be applied to accelerators and beamlines, however, like many domain-specific systems, they rely on custom implementations that are difficult to scale. Our framework builds on these efforts with standardized capability integration, scalable orchestration, and production-ready reliability across both laboratory and facility contexts, as detailed in Section III.4.
III The Alpha Berkeley Framework
The Alpha Berkeley Framework [8] is designed to move beyond proof-of-concept agents and provide a scalable, production-ready architecture for domain-adaptable workflows. The conceptual sketch is shown in Figure 1. Its distinguishing features are a classification-based approach that keeps tool usage efficient even with large capability inventories, upfront execution planning with optional human oversight, and conversational awareness that integrates multi-turn context with external memory and databases. At the same time, the framework is engineered for deployment rather than demonstration: it offers modular execution services, checkpointing and recovery for reliability, and multiple interface options ranging from a command line to browser-based GUIs and containerized services. Together, these elements make the framework both adaptable to new use cases and robust enough for sustained real-world operation in high stakes scientific environments.
@@ III.1 Task Extraction: From Conversation to Actionable Tasks
Extraction is further strengthened through multi-source data integration, combining personalized history from user memory systems, domain-specific knowledge bases for terminology resolution, and external resources such as databases, APIs, and file repositories. During this stage, the system also detects inter-task dependencies, enabling coherent sequencing in downstream workflows.
@@ III.3 Execution Planning: Upfront Orchestration
Standard ReAct-style agents invoke tools reactively, reasoning and acting in short cycles. While effective for simple tasks, this approach can become brittle in complex workflows involving multiple capabilities. Without a global view of the process, such agents risk losing track of overarching objectives, calling tools in suboptimal sequences, or overlooking critical dependencies. Errors in early steps can propagate unchecked, redundant tool calls can waste resources, and there are few natural points for human oversight before costly or irreversible actions are taken.
Our framework addresses these issues with a plan-first architecture, generating a complete execution plan before any tool is called. Multi-step workflows are coordinated with explicit input–output dependencies through context-key mapping, and planning intelligence is separated from execution logic. Plans can be serialized for inspection, modification, and resumption, with an optional planning mode that presents the plan to the user for approval or adjustment before execution begins.
@@ III.4 Execution: Production-Ready Reliability
Operating in high-stakes domains requires systems that go beyond basic orchestration, combining resilience to failures with clear operator control and flexible deployment. The framework achieves this by building on LangGraph’s checkpointing for state continuity, while adding structured error classification and recovery strategies. Errors are handled through bounded retries, re-planning, or reclassification, allowing execution to adapt gracefully without sacrificing reliability.
@@
Human-in-the-loop approval is built in as a first-class capability rather than an afterthought. Structured interrupts allow operators to inspect and approve plans, code, or memory operations before side effects occur, with all state preserved for resumption. This ensures that sensitive operations remain auditable and controllable, while still benefiting from automation when oversight is not required.
@@
The execution layer also emphasizes scientific usability. Python services support seamless switching between local and containerized execution, making it straightforward to adapt pipelines from development laptops to HPC clusters. Generated code is automatically packaged into Jupyter notebooks for inspection and editing, while artifact management captures intermediate and final outputs. Together with multiple interface options, command-line tools, browser-based workflows, and containerized deployments, these features make the framework both reliable for production use and practical for scientific environments.
@@ IV Case Studies
We illustrate the Alpha Berkeley Framework through two complementary case studies. The first is a wind farm monitoring example, presented as a illustrative tutorial example to demonstrate how the framework decomposes natural language requests into structured workflows. The second is a real deployment at the ALS, a U.S. Department of Energy synchrotron facility, where the framework is integrated directly into the accelerator control room. Together, these examples highlight both the general applicability of the framework and its readiness for production use in safety-critical scientific environments.

### v2

# Osprey: A Scalable Framework for the Orchestration of Agentic Systems

@@ Abstract
Coordinating workflows across complex systems remains a central challenge in safety-critical environments such as scientific facilities. Language-model–driven agents offer a natural interface for these tasks, but existing approaches often lack scalability, reliability, and human oversight. We introduce the Osprey Framework, a domain-agnostic, production-ready architecture for scalable agentic systems that integrate conversational context with robust tool orchestration across safety-critical domains. Our framework provides: (i) dynamic capability classification to select only relevant tools; (ii) plan-first orchestration with explicit dependencies and optional human approval; (iii) context-aware task extraction that combines dialogue history with external memory and domain resources; and (iv) production-ready execution with checkpointing, artifact management, and modular deployment. We demonstrate its versatility through two case studies: a deployment at the Advanced Light Source particle accelerator and a tutorial-style wind farm monitoring example. These results establish Osprey as a reliable and transparent framework for agentic systems across diverse high-stakes domains.
preprint: AIP/123-QED
Figure 1: Overview of the Osprey Framework workflow. First, multi-turn conversational inputs and external data sources are transformed into a structured task description. The system then dynamically identifies the relevant capabilities at each step and passes the selected tools to the execution planner. The planner generates a complete, inspectable execution plan with explicit dependencies, which is then carried out by the agent with context tracking and artifact management. The capabilities illustrated represent those from the ALS Accelerator Assistant deployment; the framework architecture supports arbitrary domain-specific capability sets.
@@ I Introduction
Coordinating large-scale, safety-critical systems, such as smart electrical grids Rajaperumal and Columbus (2025), industrial chemical processes Boskabadi et al. (2025), and scientific facilities such as particle accelerators Sulc et al. (2024), remains a central challenge in modern engineering. These environments integrate legacy and modern control systems that generate vast amounts of diverse data. Operators must act on this information under tight time constraints (for example, during accelerator fault recovery or power-grid disturbances) Yigit et al. (2025) yet retrieving, interpreting, and coordinating across systems remains cumbersome. Critical expertise is often preserved only as tacit human experience, scattered across institutional memory rather than codified in software. This tension between the need for flexibility and the rigidity of existing software architectures limits the ability to respond effectively to evolving operational demands and unexpected events.
Agentic AI Sapkota et al. (2025), Language Model (LM) powered agentic systems that orchestrate multi-step workflows through tool use, memory, and planning, offers a compelling opportunity to address these challenges. By providing a natural interface to data sources and control systems, LMs allow operators to express intent in high-level terms while delegating the orchestration of complex workflows to agents. In familiar domains such as web search Song et al. (2025) or travel planning Chen et al. (2024), LMs already demonstrate this capability, as their training reflects widespread usage patterns and benchmark tasks. By contrast, in highly specialized domains such as particle accelerators, materials science beamlines, or other expert-driven facilities, much of the relevant expertise is absent from training corpora and cannot be inferred from general usage data. Instead, the knowledge is fragmented, highly context-dependent, and often confined to experts. Deploying agentic AI in such domains requires architectures that both integrate tools effectively and translate human expertise into operational form, while preserving oversight to safeguard safety and trust.
In this paper, we introduce the Osprey Framework Osprey Developers (2025), a domain-agnostic, production-ready architecture for building scalable agentic systems across safety-critical environments. Built on LangGraph LangGraph developers (2025) for its underlying graph-based execution substrate, our framework is distinguished by four characteristics:
@@
  • Production-ready adaptability: Designed for real deployments, the framework provides modular execution environments, artifact management, checkpointing, and diverse interfaces, including a command line interface (CLI) and Open WebUI UI (2023), and containerized services.
Taken together, these results advance the framework from a proof-of-concept toward a versatile system that can be adapted to diverse safety-critical domains and deployed reliably in production environments. We illustrate the framework’s domain-agnostic capabilities with two examples: a tutorial-style wind farm monitoring scenario, and a production deployment in the Advanced Light Source (ALS) control room Hellert et al. (2025). These case studies demonstrate both the framework’s operational principles and its adaptability across different high-stakes environments.
@@ II.1 Agentic Systems and Tool Use
The foundations of agentic language models and multi-agent systems build on the classic definition of agents as entities that perceive their environment and act upon it, with rational agents aiming to maximize performance measures Russell and Norvig (2020). Extending this principle, multi-agent systems introduce shared environments in which agents may cooperate or compete. Early frameworks such as ReAct unified reasoning and acting in iterative loops for tool-based tasks Yao et al. (2023), as AutoGen Wu et al. (2023) extended this paradigm toward multi-agent conversation orchestration, enabling specialized agents to collaborate on complex objectives.
Scaling tool use in LMs has since become a central concern. Toolformer demonstrated that LMs can be trained to make self-supervised API calls Schick et al. (2023), while more recent work such as RAG-MCP addresses prompt bloat in tool selection by leveraging retrieval-augmented generation to select only relevant capabilities at inference time Gan and Sun (2025). Beyond single-turn execution, memory and retrieval play an equally important role. Retrieval-Augmented Generation (RAG) integrates external knowledge into LM reasoning Lewis et al. (2021), but remains subject to context window limitations, such as the “Lost in the Middle” phenomenon Liu et al. (2023). To overcome this, memory-augmented agents like MemGPT Packer et al. (2023) and subsequent surveys Wang et al. (2024) have highlighted methods for extending long-context reasoning. Other work formalizes agent memory as a first-class capability, emphasizing the ability to retain past interactions to inform future steps Sapkota et al. (2025). Recent examples include SciBORG Muhoberac et al. (2025), a modular agentic framework for scientific tasks that employs FSA-based memory and context awareness to provide reliable execution and interpretable state transitions. Parallel efforts in task-oriented dialogue also offer complementary insights into grounded task extraction and intent modeling Qin et al. (2023).
@@ II.2 Planning and Orchestration
Planning and orchestration in AI systems trace back to classical approaches such as STRIPS Fikes and Nilsson (1971), PDDL McDermott et al. (1998), and hierarchical task network (HTN) planning Erol et al. (1994). Recent LM-based approaches build on these ideas by incorporating structured reasoning directly into language models, exemplified by Chain-of-Thought prompting Wei et al. (2023) and Pre-Act, which combines multi-step planning with explicit reasoning steps Rawat et al. (2025). DSPy further systematizes this direction, offering a modular framework for program synthesis and prompt optimization that reduces reliance on brittle, handcrafted prompt engineering Khattab et al. (2024, 2023).
Complementing symbolic planning, graph-based architectures have emerged as powerful orchestration substrates for agents. LangGraph provides a task graph engine that enables explicit modeling of execution dependencies LangGraph developers (2025), while frameworks like PydanticAI contribute schema validation for structured outputs PydanticAI Developers (2025). Related libraries, such as Instructor Liu and Contributors (2024) and Outlines Willard and Louf (2023), offer constrained generation techniques that increase reliability in tool invocation.
Safety and human oversight remain critical in applying these systems to scientific and industrial domains. Reviews of self-driving laboratories emphasize that autonomy must integrate human intelligence for steering and safety Hysmith (2024), while approaches such as human-in-the-loop Bayesian experiment planning show how structured oversight can be embedded into experimental design Adams et al. (2024). More broadly, critical domains such as energy systems, particle accelerators, and laboratories require workflows with check-pointing, explicit approval points and safety boundaries.
@@ II.3 Domain-Specific Applications
Agentic AI has seen rapid adoption in scientific and industrial applications, where domain-specific requirements demand specialized integrations. In the natural sciences, autonomous laboratories have been identified as a key driver of accelerated discovery Szymanski et al. (2023), supported by robotic labs and LM-driven pipelines for materials exploration Vriza et al. (2025). Within imaging sciences, the PEAR framework has demonstrated multi-agent LM automation for ptychography workflows Yin et al. (2024), while other efforts have highlighted emergent autonomous research capabilities of LMs across chemistry and materials science Boiko et al. (2023a). Several domain-focused systems illustrate this diversity: ChemCrow augments LMs with chemistry-specific tools Bran et al. (2024), Co-scientist enables autonomous experimental planning in chemistry Boiko et al. (2023b), and CRISPR-GPT applies agentic orchestration to gene-editing workflows Qu et al. (2025).
Among large-scale scientific facilities, agentic approaches have also been explored for particle accelerators and synchrotron beamlines. GAIA demonstrated an early accelerator operations assistant combining retrieval, scripting, and control Mayet (2024), while VISION developed a modular AI assistant for beamline experiments, focusing on natural human–instrument interaction Mathur et al. (2025). GAIA and VISION illustrate how agentic approaches can be applied to accelerators and beamlines, however, like many domain-specific systems, they rely on custom implementations that are difficult to scale. Our framework builds on these efforts with standardized capability integration, scalable orchestration, and production-ready reliability across both laboratory and facility contexts, as detailed in Section III.4.
III The Osprey Framework
The Osprey Framework Osprey Developers (2025) is designed to move beyond proof-of-concept agents and provide a scalable, production-ready architecture for domain-agnostic workflows across safety-critical systems. The conceptual sketch is shown in Figure 1. Its distinguishing features are a classification-based approach that keeps tool usage efficient even with large capability inventories, upfront execution planning with optional human oversight, and conversational awareness that integrates multi-turn context with external memory and databases. At the same time, the framework is engineered for deployment rather than demonstration: it offers modular execution environments, checkpointing and recovery for reliability, and multiple interface options ranging from a command line to browser-based GUIs and containerized services. Together, these elements make the framework both adaptable across diverse domains and robust enough for sustained real-world operation in high-stakes scientific environments. Comprehensive documentation, including implementation guides and API references, is available online Osprey Developers (2025).
@@ III.1 Task Extraction: From Conversation to Actionable Tasks
Extraction is further strengthened through multi-source data integration, combining personalized history from user memory, domain-specific knowledge bases for terminology resolution, and external resources such as databases, APIs, and file repositories. During this stage, the system also detects inter-task dependencies, enabling coherent sequencing in downstream workflows.
@@ III.3 Execution Planning: Upfront Orchestration
Standard ReAct-style agents invoke tools reactively, reasoning and acting in short cycles. Although effective for simple tasks, this approach can become brittle in complex workflows involving multiple capabilities. Without a global view of the process, such agents risk losing track of overarching objectives, calling tools in suboptimal sequences, or overlooking critical dependencies. Errors in early steps can propagate unchecked, redundant tool calls can waste resources, and there are few natural points for human oversight before costly or irreversible actions are taken.
Our framework addresses these issues with a plan-first architecture that generates a complete execution plan before any tool is called. By check-pointing and serializing these plans, operators can review and understand what the system intends to do: inspecting, modifying, or rejecting proposed actions before any changes are made to the accelerator environment. Multi-step workflows are coordinated through explicit input–output dependencies via context-key mapping, while planning intelligence remains cleanly separated from execution logic.
In practice, the planning mode exposes the entire stepwise plan through LangGraph interrupts, giving users the ability to approve, edit, or reject individual steps prior to execution. This inspect-and-approve workflow is interface-agnostic and can be toggled at runtime with lightweight slash commands parsed by the framework gateway, for example, /planning:on to enable plan approval.
@@ III.4 Execution: Production-Ready Reliability
Operating in high-stakes domains requires systems that go beyond basic orchestration, combining resilience to failures with clear operator control and flexible deployment. The framework achieves this by building on LangGraph’s check-pointing for state continuity, while adding structured error classification and recovery strategies. Errors are handled through bounded retries, re-planning, or reclassification, allowing execution to adapt without sacrificing reliability.
@@
Human-in-the-loop approval is built in as a first-class capability rather than an afterthought. When enabled, the system pauses execution before any side effects occur and presents the current plan, code, or memory operation as a structured JSON object that can be inspected and modified. Importantly, operators can then approve, edit, or reject these proposed actions, with the full system state serialized for later resumption. This configurable mechanism keeps sensitive operations transparent, while still allowing fully automated execution when oversight is not required.
@@
The execution layer also emphasizes scientific usability. Python services support seamless switching between local and containerized execution, making it straightforward to adapt pipelines from development laptops to HPC clusters. The generated code is automatically packaged into Jupyter notebooks for inspection and editing, while artifact management captures intermediate and final outputs. Together, these features make the framework both reliable for production use and practical for scientific environments.
During execution, the framework reports structured progress and records artifacts, outputs produced by individual capabilities, such as figures, notebooks, generated code, or intermediate data passed between agents. Visual artifacts, including figures and notebooks, are rendered natively in Open WebUI, while all artifacts are registered as canonical paths in the CLI to enable inspection, reuse, and provenance tracking. Safety is enforced through explicit modes: read-only analyzes run without side effects, whereas write-enabled operations that interact with control systems (EPICS Dalesio et al. (1994) in the current deployment, extensible to other stacks) require explicit human approval under policy. The runtime supports concurrent, long-lived sessions through thread identifiers and checkpoint namespaces, allowing workflows to pause for approval and resume deterministically without loss of state, even under multi-user load.
III.5 Deployment and Operational Readiness
Beyond orchestration, the framework is engineered for production deployment with configuration-driven service management and interface consistency. A containerized stack brings up core services like pipelines, Python execution environments, and a web interface via declarative configuration, enabling reproducible development and operations. Rootless containerization and strict network boundaries reduce operational risk while preserving portability across hosts and clusters.
User interaction is unified through a gateway that normalizes input across interfaces. Both the command-line interface and Open WebUI route messages through this gateway, ensuring consistent behavior for planning mode, approval prompts, and resumable execution. Lightweight slash commands, such as /planning:on to enable plan review or /approval:selective to configure human-approval settings at runtime, allow operators to adjust oversight and diagnostics without modifying deployment settings.
Operational transparency is achieved through the structured tracking of all outputs and intermediate results. Every artifact produced by the framework can be inspected and traced to its origin, providing full provenance across interfaces. The execution layer connects seamlessly to both local and remote model providers, allowing operators to balance latency and capability while keeping all model endpoints auditable within the deployment.
Finally, the framework emphasizes modularity and clear organization rather than ad hoc expansion. Application-level prompt providers supply domain-specific instructions for orchestration, task extraction, classification, and response through a defined interface, allowing customization without modifying the underlying framework or its codebase. This modular design enables each framework deployment to incorporate its own domain-specific capabilities without changing the shared architecture. A structured component registry automatically collects available modules, such as capabilities, context handlers, data sources, and services, and records how they depend on each other. These registries can be inspected or exported for offline review, ensuring transparency and reproducibility. Together, these design choices ensure that the same agentic architecture that plans and executes complex workflows can also be configured, monitored, and maintained as a stable production system.
@@ IV Case Studies
We illustrate the Osprey Framework through two complementary case studies. The first is a wind-farm monitoring example that serves as a tutorial, demonstrating how the framework decomposes natural-language requests into structured workflows; a comprehensive, step-by-step version of this tutorial is available in the project’s code documentation on GitHub Osprey Developers (2025). The second is a production validation conducted at the Advanced Light Source (ALS) Hellert et al. (2024), a U.S. Department of Energy synchrotron facility, where the framework underlies the Accelerator Assistant system Hellert et al. (2025). Together, these examples highlight both the general applicability of the framework and its readiness for deployment in safety-critical scientific environments.